From nch@roe.ac.uk Mon Apr 11 13:03:03 2011 Date: Fri, 8 Apr 2011 16:47:28 +0100 (BST) From: Nigel Hambly To: WFCAM Science Archive Team -- Eckhard Sutorius , Mike Read , Mark Holliman , Nigel Hambly , Nicholas Cross , Rob Blake , Ross Collins Cc: CCs for WSA weekly meeting minutes distribution -- Andrew Lawrence , Jim Emerson , "Noddle, Keith -- Keith Noddle" , Keith Noddle , Lorenzo Rimoldini , Mike Irwin , Bob Mann , Stephen Warren , Tom Kerr Subject: WFAU VDFS Science Archive project meeting minutes, 1/4/11 Minutes of WFAU VDFS Science Archive meeting: April 1st 2011 ------------------------------------------------------------------------- ------------------------------------------------------------------------- Present: NCH, ETWS, MAR, NJC, RPB, MSH, RGM Apologies: JPE, KTN, AL, RSC DoNM: Friday April 15th 2011 in the Vista Hut Actions discharged this week: ----------------------------- ACTION: RPB to investigate LTO4 tape library backups from the linux client side. Discharged; NCH asked RPB to make a new backup of the legacy SuperCOS datasets now that this facility is available (existing LTO-3 backup now more than 3yrs old). ACTION: MSH and KTN to spec/tender for another cluster node and another web server Discharged ACTION: RPB to make the full-blown UKIDSS DR5 including UDS/GPS world public. Discharged Actions partly discharged but continuing: ----------------------------------------- ACTION: NCH to ask RPB to look into the perennial functions permissions problem. RPB to communicate with RSC to insert into CU19 (ACTION RSC) Actions carried forward from 11/03/11 meeting --------------------------------------------- ACTION: RPB to look into automated RAID monitoring across all servers. Continues; ramses9 (most recent SW installation) has such a facility; it was noted that on the linux side monitoring is working as normal. ACTION: NCH to ensure a clear schedule for public release of VISTA survey DBs is agreed between all interested parties ACTION: MSH to set up ramses10 in the cluster rack and configure as a further ingest DB server. Specific points and new actions: -------------------------------- Project management: All noted the newly reconstituted WFAU quarterly meetings to commence this week on Tuesday. NCH reported back from the recent VHS Ultra-Cool Dwarfs science working group telecon, where several science exploitation issues relating to the VSA were discussed (see below). NJC reported back from the VMC survey team telecon, and it was noted that the team seems happy with the latest release products. WFCAM & VISTA updates: There has been a delay in the VISTA mirror re-coating schedule; the date when science operations can be continued is still open. Bad weather continues on Mauna Kea. Loadsa time to catch up with all the reprocessing then... Comments and issues arising from CASU minutes: No new minutes this (i.e. last!) week. Networking: ETWS noted the network continues to be a little unreliable, and suspects intermittent problems on Janet or down south. Despite this, transfers (when running) holding up at a slightly reduced rate of 20 Mbyte/s, and all v1.1 reprocessed VISTA data is now copied up to the end of Jan 2011. WSA/VSA Operations: ETWS noted: - The transfer and metadata ingests for all the v1.1 data up to 2011-01-31 into a seperate v1.1 database will finish today. - Implemented the illumination correction for VISTA. - Updated the monitor script to allow data to be read from multiple databases. Also updated and enhanced scripts to check the transfers and ingests. The team again reviewed the status of the various VISTA surveys, particularly regarding the ESO deliverables: o VMC: another candidate ESO release done for the PI... so far, so good o VHS: main new (!) requirements to emerge from a recent telecon are monthly uploads to ESO beginning with P84/dry-run, and incorporation of the imminent WISE catalogue release into the VSA - http://wise2.ipac.caltech.edu/docs/release/prelim/preview.html o VVV: no goalpost shifts this week o VIKING: v1.0 tiling should finish next (this) week and then a new release DB will be created. The possibility of a 25% v1.1 release DB by the end of April is being considered in consultation with the PI. o VIDEO: still unclear as to requirements for "band-merged products" (whatever that means) for the current deadline. o UltraVISTA: still SEP. Hardware and Systems: MSH noted that the old server room C1 is being cleared, and that he has checked that there is nothing valuable lying around in there belonging to WFAU. MSH also noted that shepseskaf is now on the SRIF network. Software: RSC noted: "Implemented in our ESO release script changes to the official ESO-SAF metadata specification for single/multi-band catalogues and some additional changes following feedback from PIs about the metadata." NJC has fixed some problems with SExtractor running as part of the DB-driven products pipeline, and has sorted out the VIKING tiling problems (fingers crossed). Survey Data Release: No news this week. Non-survey Data Release: No news this week. Astrogrid deployment & Data Analysis services: No news this week. Miscellaneous: Nowt else this week (resisting all temptation to put something daft in given the date of these minutes... mind you, don't expect anybody reads this far. Dunno why I bother.) ============================================================= Nigel Hambly Tel: +44-131-668-8234 Institute for Astronomy Fax: +44-131-668-8416 School of Physics and Astronomy University of Edinburgh Email: nch@roe.ac.uk Royal Observatory Blackford Hill Edinburgh EH9 3HJ The University of Edinburgh is a charitable body, registered in Scotland, with registration number SC005336. ============================================================= -- Scanned by iCritical.